OpenStudy (anonymous):

please help i dont get it? For your birthday you are given $150. After three weeks you have $120. Two weeks after that, you have $100. If the amount that you spend each week remains constant, how much will you have 10 weeks after your birthday? (1 point) $40 $50 $60 $70

OpenStudy (anas.p):

In 3 weeks you are spending $30 In 2 weeks you are spending $20 So you are spending $10 a week You received $150 for your birthday. 10 weeks later you will have spend $100 ($10*10) So you will have $150-$100 left that is $50.

OpenStudy (anonymous):

this is stupid but 50-10=40 so how are you only spending 10 a week

OpenStudy (anas.p):

150 after 3 weeks you have 120 So 150-120=30 30 in 3 weeks 30/3=1 Again From 120 you have 100 left after 2 weeks So 120-100=20 20 in 2 weeks 20/2=10

OpenStudy (anonymous):

ok i see thanks!